OpenStudy (anonymous):

How much grapefruit juice is needed to make 16 servings of fruit punch? A. 4 c B. 6 c C. 7 c D. 8 c

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

do you have a function or equations?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

idk how much grape juice is it per serving.

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

6 looks liike you just double it

OpenStudy (anonymous):

i think it's A. am i right

OpenStudy (anonymous):

B :) is 6 you just double it

OpenStudy (anonymous):

6+6=12

OpenStudy (anonymous):

16/8=2 8/4=2 2/1=2 all accross the board of your pic you get 2 :) so just go like this x/3=2 x=3*2 multiply both sides by 2 and you get 6